I am 32 years old, my nephew is 7. At what age will we both be, when I am exactly 3 times his age? Thank you for your help
Answer by fractalier(6550)

Call x the number of years in the future...then we have
32 + x = 3(7 + x)
32 + x = 21 + 3x
11 = 2x
x = 5.5 years
Thus you will be 37.5 and he will be 12.5 years old.